OpenStudy (anonymous):

The molar mass of magnesium (Mg) is 24.30 g/mol. There are 6.02 1023 atoms in 1 mol. How many atoms are present in 48.60 g of Mg?

OpenStudy (accessdenied):

We are given the mass of Mg. We'll need to first convert it to moles using molar mass (g/mol). Then we can convert to number of atoms as there are \(6.02 \times 10^{23} \) atoms per mole.

OpenStudy (accessdenied):

If you work well with units, it is usually very clear what step must be done. We start with grams. To get to moles, we need to divide by (grams/mole), molar mass. Then to get to atoms, we multiply by atoms per mole. \( \cancel{\text{grams}} \ / \ \left( \dfrac{\cancel{\text{grams}}}{\text{moles}} \right) = \text{moles} \) \( \cancel{\text{moles}} \times \dfrac{\text{atoms}}{\cancel{\text{mole}}} = atoms \) These are the unit conversions. Now you just need to compare it to your information and apply the process!